A fuel distributor has 161,798 gallons of fuel with 1% sulfur content, which exceeds pollution control standards of 0.6% sulfur content. How many gallons of fuel oil with a 0.1% sulfur content must be added to the 161,798 gallons to obtain fuel oil that complies with the pollution control standards? Round up to the next higher gallon.

Answer by stanbon(75887)

A fuel distributor has 161,798 gallons of fuel with 1% sulfur content, which exceeds pollution control standards of 0.6% sulfur content. How many gallons of fuel oil with a 0.1% sulfur content must be added to the 161,798 gallons to obtain fuel oil that complies with the pollution control standards? Round up to the next higher gallon.
-----------------------------
1% oil DATA:
Amt = 161,798 gals ; Amt of sulfur = 0.01*161,798 = 1617.98 gallons
---------------------
0.1% oil DATA:
Amt = x gals ; Amt of sulfur = 0.001x gallons
----------------------
0.6% Mixture DATA:
Amt = 161798+x ; Amt of sulfur = 0.006(161798-x) = 970.788 - 0.006x gals
-----------------------
EQUATION:
sulfur + sulfur = sulfur
1617.98 + 0.001x = 970.788 - 0.006x
647.192 = 0.005x
x = 129,438.4 gallons of 0.1% oil must be added to bring the tank down
to 0.6% sulfur content.
==============
Cheers,
Stan H.